# A framework for the valuation of insurance liabilities by production cost

### v1

@@ 3.1 General
We denote by $\mathbb{R}^{d+1}_{\geq 0}$ for $d\in\{0,1,\ldots\}$ the set of $x=(x_{0},\ldots,x_{d})\in\mathbb{R}^{d+1}$ with $x_{k}\geq 0$ for all $k\in\{0,1,\ldots,d\}$. For $a,b\in\mathbb{R}$, we use $(a)_{+}=\max\{0,a\}$ and $a\wedge b=\min\{a,b\}$. For a subset $A\in\mathcal{F}$, the indicator function $1_{A}$ is $11_{A}(\omega)=$ 1 for $\omega\in A$ and $01_{A}(\omega)=$ 0 otherwise. We write “iff” for “if and only if”.
Cash flows. We distinguish non-negative incoming cash flows, generally denoted by the letter $Z_{t}$, typically from assets, from non-negative outgoing cash flows $X_{t}$, typically from liabilities, at dates $t\in D[0,T]$. Incoming and outgoing is often in relation to the insurance company under consideration but also sometimes to an investment strategy (Definition 3.1). We distinguish contractually specified non-negative cash flows $Z_{t}$ and $X_{t}$ from non-negative actual cash flows $\tilde{Z}_{t}$ and $\tilde{X}_{t}$ that consider default. Incoming cash flows from selling and outgoing cash flows from buying an asset are not considered as cash flows “of” the asset.
Tradables. We consider a financial market with $d+1$ tradables defined as financial asset instruments enumerated by $k=0,1,\dots,d$ with $\mathcal{F}_{t}$-measurable non-negative reliable market prices $S_{t}=(S_{t}^{0},S_{t}^{1},\ldots,S_{t}^{d})$ in $\mathbb{R}^{d+1}_{\geq 0}$ at dates $t\in D[0,T]$. We make in general no further assumptions specifically about the tradable $k=0$. A market price is reliable (as an idealization) if the instrument is “fully liquid”, i.e. any quantity can be bought or sold instantly for cash equal to the market price (disregarding transaction costs and other market frictions). For the company holding them as assets, tradables may generate $\mathcal{F}_{t}$-measurable incoming actual cash flows $\tilde{Z}_{t}^{k}\geq 0$ ($k\in\{0,\dots,d\}$) such as coupons for bonds and dividends for equities, for simplicity only at dates $t\in D[0,T]$. As a further simplification, we assume that the tradables have no outgoing cash flows. We use the convention that the market price $S_{t}^{k}$ at a cash flow date $t\in D[0,T]$ does not include, i.e. is “immediately after” the cash flow $\tilde{Z}_{t}^{k}$. Because cash flows are relevant in the framework for default and for illiquid assets, we use as $S_{t}^{k}$ the actually observed market prices and not total return indices derived by assuming that incoming cash flows are automatically reinvested in the tradable.
In view of the asymmetry between assets and liabilities, we first consider in Section 3.2 only non-negative (investment) strategies of tradables (i.e. assets), that is, with only non-negative units of tradables. In Section 3.3, we then define general strategies including short positions in tradables.
@@ 3.2 Non-negative strategies, illiquid assets
Strategies for producing insurance liabilities in general include incoming and outgoing cash flows $\tilde{Z}_{t}$ and ${X}_{t}$ from the insurance liabilities such as premiums and claims payments, respectively. As incoming and outgoing cash flows do not in general cancel each other out, we consider non-negative strategies $\phi$ of tradables with finite discrete trading dates $t\in D[t_{min},t_{max}]$ that are not necessarily self-financing. A non-negative (asset) portfolio $\phi_{t_{j+1}}\geq 0$ of tradables for $t_{j}\in D[t_{min},t_{max}]$ with $t_{j}<t_{max}$ is defined to be an $\mathcal{F}_{t_{j}}$-measurable random vector on $\mathbb{R}^{d+1}_{\geq 0}$. In particular, for $t\equiv t_{j+1}$, the number $\phi_{t}^{k}$ of units of any tradable $k$ is non-negative. The incoming (actual) cash flows of the portfolio $\phi_{t}$ are $\tilde{Z}_{t}^{\phi_{t}}=\sum_{k=0}^{d}\phi_{t}^{k}\cdot\tilde{Z}_{t}^{k}$, and we write $\phi_{t}\cdot S_{s}=\sum_{k=0}^{d}\phi_{t}^{k}\cdot S_{s}^{k}$.
@@
For $D\equiv D[t_{min},t_{max}]$, a non-negative (investment) strategy $\phi=(\phi_{t})_{t\in D}$ of tradables with cash flows $(\tilde{Z}_{t})_{t\in D}$ incoming to and cash flows $({X}_{t})_{t\in D}$ outgoing of the strategy is a predictable sequence of non-negative portfolios $\phi_{t_{j}}$ for ${t_{j}}\in D\cup\{\gamma(t_{max})\}$, i.e. $\phi_{t_{min}}$ is $\mathcal{F}_{t_{min}}$-measurable and $\phi_{t_{j}}$ is $\mathcal{F}_{t_{j-1}}$-measurable for $t_{j}>t_{min}$ in $D\cup\{\gamma(t_{max})\}$, such that at any date $t_{j}\in D$,
@@
where we define $\tilde{Z}_{t_{j}}^{\phi}=\tilde{Z}_{t_{j}}^{\phi_{t_{j}}}$. The strategy $\phi$ is self-financing iff $\tilde{Z}_{t_{j}}=X_{t_{j}}$ for any $t_{j}\in D$, i.e.
$\phi_{t_{j+1}}\cdot S_{t_{j}}=\phi_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\phi}$ (3)
The value $v_{t_{j}}(\phi)$ at $t_{j}\in D$ is defined to be the market price of the portfolio $\phi_{t_{j+1}}$ after the cash flows at date $t_{j}\in D$,
@@
Note that (2) does not determine a strategy $\phi$.
As an illustration, consider units of only one tradable $k\in\{1,\ldots,d\}$ with non-zero incoming cash flows $\tilde{Z}_{t}^{\phi^{k}}\neq 0$. The strategy to invest the incoming cash flows $\tilde{Z}_{t}^{\phi^{k}}$ in additional units of the tradable $k$ has cash flows $\tilde{Z}_{t}={X}_{t}=0$ incoming to and outgoing of the strategy, respectively, and is thus self-financing according to Definition 3.1. In contrast, the strategy of passively holding the units unchanged has $\tilde{Z}_{t}=0$ and ${X}_{t}=\tilde{Z}_{t}^{\phi^{k}}\neq 0$ and so is not self-financing. Here, the cash flows ${X}_{t}$ are outgoing relative to the strategy but may remain as cash within the company.
The following property of “consistency” of a set of tradables takes in our framework a role similar to the classical no-arbitrage condition.
@@
This is a condition on tradables because the strategy is not changed between the two dates $t$ and $\gamma(t)$. The corresponding property for strategies is the following: if, at some date $t_{j}\in D$, the values of the two strategies $\phi$ and $\theta$ satisfy $v_{t_{j}}(\phi)\geq v_{t_{j}}(\theta)$, then $v_{t_{\ell}}(\phi)\geq v_{t_{\ell}}(\theta)$ at any earlier date $t_{\ell}\in D$, $t_{\ell}<t_{j}$. This property depends on consistency of tradables as well as on the conversion of the strategies according to (2) at the intermediate dates, specifically (6) below. To show the property, we start from two non-negative strategies $\phi$ and $\theta$ with cash flows $\tilde{Z}_{t},X_{t}$ and $\tilde{Z}_{t}^{\prime},X_{t}^{\prime}$, respectively, such that $v_{t_{j}}(\phi)\geq v_{t_{j}}(\theta)$. Using the definition (4) of the value and (2), this implies $\phi_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\phi}+\tilde{Z}_{t_{j}}-X_{t_{j}}\geq\theta_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\theta}+\tilde{Z}_{t_{j}}^{\prime}-X_{t_{j}}^{\prime}$. If $\phi$ and $\theta$ have the same “net cash flows”, i.e.
@@
Illiquid assets. In addition to tradables, we consider a non-negative portfolio $\psi=(\psi_{0}^{1},\ldots,\psi_{0}^{l})\in\mathbb{R}^{l}_{\geq 0}$ of a finite number of “fully” illiquid assets at date $t=0$, making the extreme assumption that these cannot be sold or bought, so $\psi$ is “held to maturity” (HTM). We assume that any illiquid asset $\ell$ has $\mathcal{F}_{t}$-measurable actual incoming cash flows $\tilde{Z}_{t}^{\ell}\geq 0$ and no outgoing cash flows. The corresponding non-negative strategy also denoted by $\psi$ has cash flows $\tilde{Z}_{t}^{\psi}$ at dates $t\in D[0,T]$ given by $\tilde{Z}_{t}^{\psi}=\sum_{\ell=1}^{l}\psi_{0}^{\ell}\cdot\tilde{Z}_{t}^{\ell}$. This corresponds to a (not self-financing) investment strategy as in Definition 3.1 with no incoming cash flows $\tilde{Z}_{t}=0$ and outgoing cash flows $X_{t}=\tilde{Z}_{t}^{\psi}$, i.e. with $\psi_{\gamma(t)}\cdot S_{t}=\psi_{t}\cdot S_{t}$ corresponding to (2).
@@ 3.3 Short positions, general strategies, restrictions
Short position. For positive units $\phi^{k}>0$ of the tradable $k$ as an asset, the corresponding negative units $\theta^{k}=-\phi^{k}<0$ by “going short” are a liability. The asset that is the “counterposition” to this liability in general is different from the original asset. E.g. “going short” in cash, i.e. borrowing cash by some debt contract produces an asset position of the cash lender in the debt contract that is different from simply holding cash. This illustrates that asset and liability from “going short” need to be defined through an agreement between the counterparties, with per-se several possible definitions. (Further, the agreement is exposed to failure of the liability counterparty.) In the definition we choose, the original asset is in particular ”replicated” in the following sense: the asset counterparty of the agreement receives the incoming actual (not the contractually specified) cash flows $\phi^{k}\cdot\tilde{Z}_{t}^{k}$ of the units of the tradable $k$ (e.g. bond coupons, dividends) and in addition has the option at any time to get the current market price $\phi^{k}\cdot S_{t}^{k}$ and any incoming cash flows $\phi^{k}\cdot\tilde{Z}_{t}^{k}$ of the position in cash, which terminates the agreement. To use negative units of tradables in production, we need to assume that such agreements can be set up. We may further assume that they can be closed out.
We define short positions directly as a liability $\mathcal{L}(\phi)$ for non-negative strategies $\phi$ as in Definition 3.1 with incoming cash flows $\tilde{Z}_{t}=0$ and cash flows $X_{t}$ outgoing of the strategy $\phi$ but that are assumed to initially remain with the company. We consider general cash flows $X_{t}$, with the incoming cash flows $\tilde{Z}_{t}^{\phi}$ going out being a typical special case. We define $\mathcal{L}(\phi)$ as a “replication” of the strategy $\phi$ to the corresponding asset holder, in particular with the cash flows $X_{t}$ paid out to the asset holder, with an option for the asset holder to terminate the agreement and get the value in cash. Short positions in a tradable are a special case.
@@
Let $\phi\geq 0$ be a non-negative strategy on $D\equiv D[t_{min},t_{max}]$ with outgoing cash flows $X_{t}$ and no incoming cash flows $\tilde{Z}_{t}=0$. The short position liability $\mathcal{L}({\phi})$ on $D$ is defined using a stopping time $\tau\in D$ by incoming cash flows $\tilde{Z}_{t}^{\mathcal{L}(\phi)}=0$ ($t\in D$) and outgoing cash flows for $t\in D$,
@@
Portfolios containing negative units of a tradable $k\in\{0,1,\ldots,d\}$ are defined to be available for production on $D$ iff it is possible at any $t\in D$ for the market price to take on a short position of the tradable $k$ and available for production with close out iff it is in addition possible at any $t\in D$ for the market price to close out an existing short position.
@@
General strategy. A non-negative strategy as in Definition 3.1 is characterized by asset portfolios and the portfolio conversion equation (2). For the extension to general strategies also containing short positions, we formally decompose portfolios $\phi_{t}\in\mathbb{R}^{d+1}$ into $\phi_{t}=\phi_{t}^{+}-\phi_{t}^{-}$ with non-negative portfolios $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$ and interpret this as an asset portfolio $\phi_{t}^{+}$ together with a liability *$\mathcal{L}^{*}(\phi^{-})$ to be defined through its cash flows. Inserting $\phi_{t}=\phi_{t}^{+}-\phi_{t}^{-}$ into the formal portfolio conversion equation (2) for $\phi$ at date $t\in D$ and rearranging we get a portfolio conversion equation for the non-negative strategy $\phi^{+}$:
@@
In view of this, we define the cash flows for the liability **$\mathcal{L}^{*}\equiv\mathcal{L}^{*}(\phi^{-})$ at dates $t$ to be the outgoing cash flow $X_{t}^{\mathcal{L}^{*}}=\phi_{t}^{-}\cdot S_{t}+\tilde{Z}^{\phi^{-}}_{t}$ above from closing out at date $t$ a short position in the portfolio $\phi_{t}^{-}$ and the incoming cash flow $\tilde{Z}_{t}^{\mathcal{L}^{*}}=\phi_{\gamma(t)}^{-}\cdot S_{t}$ above from taking on a new short position in the portfolio $\phi_{\gamma(t)}^{-}$ provided that $t<t_{max}$.
@@
Assume that portfolios containing negative units of tradables are available for production with close out. A strategy $\phi=\phi^{+}-\phi^{-}$ on $D\equiv D[t_{min},t_{max}]$ with cash flows $\tilde{Z}_{t}$ incoming to and cash flows $X_{t}$ outgoing of the strategy is defined as the non-negative strategy $\phi^{+}$ with cash flows $\tilde{Z}_{t}$ and $X_{t}$ together with the liability **$\mathcal{L}^{*}\equiv\mathcal{L}^{*}(\phi^{-})$ defined by the outgoing and incoming cash flows for any $t\in D$,
@@
In certain applications, strategies may have to be restricted e.g. to investing only in a subset of tradables, i.e. $\phi_{t}^{k}=0$ for one or several $k\in\{0,1,\ldots d\}$.
@@
For a linear subspace $\mathcal{R}$ of $\mathbb{R}^{d+1}$ and a strategy $\phi$ on $D\equiv D[t_{min},t_{max}]$, we define, by slight abuse of notation, $\phi\in\mathcal{R}$ iff $\phi_{t}\in\mathcal{R}$ for any $t\in D$, ϕ+italic-$\phi\in\mathcal{R}^{+}$ iff $0\leq\phi_{t}\in\mathcal{R}$ for any $t\in D$, and ϕitalic-$\phi\in\mathcal{R}^{\prime}$ iff $\phi_{t}\in\mathcal{R}$ and $v_{t}(\phi)\geq 0$ for any $t\in D$. (Note: $\phi\in\mathcal{R}\Rightarrow\phi^{\pm}\in\mathcal{R}$.)

### v2

@@ 2.1 General
We denote by $\mathbb{R}^{d+1}_{\geq 0}$ for $d\in\{0,1,\ldots\}$ the set of $x=(x_{0},\ldots,x_{d})\in\mathbb{R}^{d+1}$ with $x_{k}\geq 0$ for all $k\in\{0,1,\ldots,d\}$. For $a,b\in\mathbb{R}$, we write $(a)_{+}=\max\{0,a\}$ and $a\wedge b=\min\{a,b\}$. For a subset $A\in\mathcal{F}$, the indicator function $1_{A}$ is $11_{A}(\omega)=$ 1 for $\omega\in A$ and $01_{A}(\omega)=$ 0 otherwise. We write “iff” for “if and only if”.
Cash flows. We distinguish non-negative cash inflows, generally denoted by the letter $Z_{t}$, typically from assets, from non-negative cash outflows $X_{t}$, typically from liabilities, at dates $t\in D[0,T]$. We refer to in- and outflows often in relation to the insurance company under consideration but also sometimes in relation to an investment strategy (Definition 2.1). We distinguish contractually specified non-negative cash flows $Z_{t}$ and $X_{t}$ from non-negative actual cash flows $\tilde{Z}_{t}$ and $\tilde{X}_{t}$ that consider failure. Cash inflows from selling and cash outflows from buying an asset are not considered as cash flows “of” the asset. We consider a cash flow simply as a “flow” of tradables (see below) with a specific value, not restricting to cash in a specific currency.
Tradables. We consider a financial market with $d+1$ tradables defined as financial asset instruments enumerated by $k=0,1,\dots,d$ with $\mathcal{F}_{t}$-measurable non-negative reliable market prices $S_{t}=(S_{t}^{0},S_{t}^{1},\ldots,S_{t}^{d})$ in $\mathbb{R}^{d+1}_{\geq 0}$ with $S_{t}\neq 0$ at dates $t\in D[0,T]$. We make in general no further assumptions, specifically about the tradable $k=0$. The market price of an instrument is reliable (as an idealization) if the instrument is “fully liquid”, i.e. any quantity can be bought or sold instantly for a cash flow equal to the market price (disregarding transaction costs and other market frictions). For the company holding them as assets, tradables may generate $\mathcal{F}_{t}$-measurable actual cash inflows $\tilde{Z}_{t}^{k}\geq 0$ ($k\in\{0,\dots,d\}$) such as coupons for bonds and dividends for equities, for simplicity only at dates $t\in D[0,T]$. As a further simplification, we assume that the tradables have no cash outflows. We use the convention that the market price $S_{t}^{k}$ at a cash flow date $t\in D[0,T]$ does not include, i.e. is “immediately after” the cash inflow $\tilde{Z}_{t}^{k}$. Because cash flows are relevant in the framework for default and for illiquid assets, we use as $S_{t}^{k}$ the actually observed market prices and not total return indices derived by assuming that cash inflows are automatically reinvested in the tradable.
In view of the asymmetry between assets and liabilities, we first consider in Section 2.2 only non-negative (investment) strategies of tradables (i.e. assets), that is, with only non-negative units of tradables. In Section 2.3, we then define general strategies that allow for negative units of tradables.
@@ 2.2 Non-negative strategies, illiquid assets
Non-negative strategies. Strategies for producing insurance liabilities in general include cash inflows $\tilde{Z}_{t}$ and outflows ${X}_{t}$ from the insurance liabilities such as premiums and claims payments, respectively, that do not in general cancel each other out. For this reason, we consider non-negative strategies $\phi$ of tradables with finite discrete trading dates $t\in D[t_{min},t_{max}]$ that are not necessarily self-financing. A non-negative (asset) portfolio $\phi_{t_{j+1}}\geq 0$ of tradables for $t_{j}\in D[t_{min},t_{max}]$ with $t_{j}<t_{max}$ is defined to be an $\mathcal{F}_{t_{j}}$-measurable random vector on $\mathbb{R}^{d+1}_{\geq 0}$. Abbreviating $t\equiv t_{j+1}$, in particular, the number $\phi_{t}^{k}$ of units of any tradable $k$ is non-negative. The (actual) cash inflows of the portfolio $\phi_{t}$ are $\tilde{Z}_{t}^{\phi_{t}}=\sum_{k=0}^{d}\phi_{t}^{k}\cdot\tilde{Z}_{t}^{k}$, and we write $\phi_{t}\cdot S_{s}=\sum_{k=0}^{d}\phi_{t}^{k}\cdot S_{s}^{k}$.
@@
For $D\equiv D[t_{min},t_{max}]$, a non-negative (investment) strategy $\phi=(\phi_{t})_{t\in D}$ of tradables with cash inflows $(\tilde{Z}_{t})_{t\in D}$ and outflows $({X}_{t})_{t\in D}$ to the strategy is a predictable sequence of non-negative portfolios $\phi_{t_{j}}$ for ${t_{j}}\in D\cup\{\gamma(t_{max})\}$, i.e. $\phi_{t_{min}}$ is $\mathcal{F}_{t_{min}}$-measurable and $\phi_{t_{j}}$ is $\mathcal{F}_{t_{j-1}}$-measurable for $t_{j}>t_{min}$ in $D\cup\{\gamma(t_{max})\}$, such that at any date $t_{j}\in D$,
@@
where we define $\tilde{Z}_{t_{j}}^{\phi}=\tilde{Z}_{t_{j}}^{\phi_{t_{j}}}$. The value $v_{t_{j}}(\phi)$ of a strategy $\phi$ at $t_{j}\in D$ is defined to be the market price of the portfolio $\phi_{t_{j+1}}$ after the cash flows at date $t_{j}\in D$,
@@
The strategy $\phi$ is self-financing iff $\tilde{Z}_{t_{j}}=X_{t_{j}}$ for any $t_{j}\in D$, i.e. (2) becomes
$\phi_{t_{j+1}}\cdot S_{t_{j}}=\phi_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\phi}$ (4)
Remark 2.2.
Note that we do not consider the cash inflows $\tilde{Z}_{t_{j}}^{\phi}$ to be part of the cash inflows $(\tilde{Z}_{t})_{t\in D}$ of the strategy. Also note that (2) in general does not determine the portfolio $\phi_{t_{j+1}}$.
As an illustration, consider units of only one tradable $k\in\{0,\ldots,d\}$ with non-zero cash inflows $\tilde{Z}_{t}^{\phi^{k}}\neq 0$. The strategy to invest the cash inflows $\tilde{Z}_{t}^{\phi^{k}}$ in additional units of the tradable $k$ has cash flows $\tilde{Z}_{t}={X}_{t}=0$ incoming to and outgoing of the strategy, respectively, and is thus self-financing according to Definition 2.1. In contrast, the strategy of passively holding the units unchanged has $\tilde{Z}_{t}=0$ and ${X}_{t}=\tilde{Z}_{t}^{\phi^{k}}\neq 0$ and so is not self-financing. Here, the cash flows ${X}_{t}$ are outgoing relative to the strategy but may remain within the company.
In our framework, the following property of “consistency” of a set of tradables takes a role similar to the classical no-arbitrage condition.
@@
This is a condition on tradables because the strategy is not changed between the two dates $t$ and $\gamma(t)$. The corresponding property for strategies is the following: if, at some date $t_{j}\in D$, the values of the two strategies $\phi$ and $\theta$ satisfy $v_{t_{j}}(\phi)\geq v_{t_{j}}(\theta)$, then $v_{t_{\ell}}(\phi)\geq v_{t_{\ell}}(\theta)$ at any earlier date $t_{\ell}\in D$ with $t_{\ell}<t_{j}$. This property depends on consistency of tradables as well as on the conversion of the strategies according to (2) at the intermediate dates, as we now show. To investigate when the property holds, consider two non-negative strategies $\phi$ and $\theta$ with cash flows $\tilde{Z}_{t},X_{t}$ and $\tilde{Z}_{t}^{\prime},X_{t}^{\prime}$, respectively, such that $v_{t_{j}}(\phi)\geq v_{t_{j}}(\theta)$. Using the definition (3) of the value and (2), this implies $\phi_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\phi}+\tilde{Z}_{t_{j}}-X_{t_{j}}\geq\theta_{t_{j}}\cdot S_{t_{j}}+\tilde{Z}_{t_{j}}^{\theta}+\tilde{Z}_{t_{j}}^{\prime}-X_{t_{j}}^{\prime}$. If we then assume that $\phi$ and $\theta$ have the same “net cash flows”, i.e.
@@
Illiquid assets. In addition to tradables, we consider a non-negative portfolio $\psi=(\psi_{0}^{1},\ldots,\psi_{0}^{l})\in\mathbb{R}^{l}_{\geq 0}$ of a finite number of “fully” illiquid assets at date $t=0$, making the extreme assumption that these cannot be sold or bought, so $\psi$ is “held to maturity” (HTM). We assume that any illiquid asset $\ell$ has $\mathcal{F}_{t}$-measurable actual cash inflows $\tilde{Z}_{t}^{\ell}\geq 0$ and no cash outflows. The corresponding non-negative strategy also denoted by $\psi$ has cash flows $\tilde{Z}_{t}^{\psi}$ at dates $t\in D[0,T]$ given by $\tilde{Z}_{t}^{\psi}=\sum_{\ell=1}^{l}\psi_{0}^{\ell}\cdot\tilde{Z}_{t}^{\ell}$. This corresponds to a (not self-financing) investment strategy as in Definition 2.1 with no cash inflows $\tilde{Z}_{t}=0$ and cash outflows $X_{t}=\tilde{Z}_{t}^{\psi}$, i.e. with $\psi_{\gamma(t)}\cdot S_{t}=\psi_{t}\cdot S_{t}$ corresponding to (2).
@@ 2.3 Short positions, general strategies, restrictions
Short position. For positive units $\phi^{k}>0$ of the tradable $k$ as an asset, the corresponding negative units $\theta^{k}=-\phi^{k}<0$ by “going short” are a liability. The asset that is the “counterposition” to this liability in general is different from the original asset. E.g. “going short” in cash, i.e. borrowing cash by some debt contract produces an asset position of the cash lender in the debt contract that is different from simply holding cash. This illustrates that asset and liability from “going short” need to be defined through an agreement between the counterparties, with per-se several possible definitions. (Further, the agreement is exposed to failure of the liability counterparty.) In the definition we choose, the original asset is in particular ”replicated” in the following sense: the asset counterparty of the agreement receives the actual (not the contractually specified) cash inflows $\phi^{k}\cdot\tilde{Z}_{t}^{k}$ of the units of the tradable $k$ (e.g. bond coupons, dividends) and in addition has the option at any time to get the current market price $\phi^{k}\cdot S_{t}^{k}$ and any cash inflows $\phi^{k}\cdot\tilde{Z}_{t}^{k}$ of the position in cash, which terminates the agreement. To use negative units of tradables in production, we need to assume that such agreements can be set up. We may further assume that they can be closed out.
We define short positions directly as a liability $\mathcal{L}(\phi)$ for non-negative strategies $\phi$ as in Definition 2.1 with cash inflows $\tilde{Z}_{t}=0$ and cash outflows $X_{t}$ to the strategy $\phi$ but that are assumed to initially remain with the company. We consider general cash outflows $X_{t}$, with the cash inflows $\tilde{Z}_{t}^{\phi}$ going out being a typical special case. We define $\mathcal{L}(\phi)$ as a “replication” of the strategy $\phi$ to the corresponding asset holder, in particular with the cash flows $X_{t}$ paid out to the asset holder, with an option for the asset holder to terminate the agreement and get the value in cash. Short positions in a tradable are a special case.
@@
Let $\phi\geq 0$ be a non-negative strategy on $D\equiv D[t_{min},t_{max}]$ with cash outflows $X_{t}$ and no cash inflows $\tilde{Z}_{t}=0$. The short position liability $\mathcal{L}({\phi})$ on $D$ is defined using a stopping time $\tau\in D$ by cash inflows $\tilde{Z}_{t}^{\mathcal{L}(\phi)}=0$ ($t\in D$) and cash outflows for $t\in D$,
@@
Portfolios containing negative units of a tradable $k\in\{0,1,\ldots,d\}$ are said to be available for production on $D$ iff it is possible at any $t\in D$, for the market price of the corresponding positive units of the tradable $k$, to borrow the units of the tradable $k$ by taking on the corresponding liability, and available for production with close out iff it is in addition possible at any $t\in D$ for the market price to close out (pay back) such an existing liability.
@@
General strategy. A non-negative strategy as in Definition 2.1 is characterized by asset portfolios and the portfolio conversion equation (2). For the extension to general strategies also containing short positions, we formally decompose portfolios $\phi_{t}\in\mathbb{R}^{d+1}$ into $\phi_{t}=\phi_{t}^{+}-\phi_{t}^{-}$ with non-negative portfolios $\phi_{t}^{\pm}=(\pm\phi_{t})_{+}\geq 0$ and interpret this as an asset portfolio $\phi_{t}^{+}$ together with a liability $\mathcal{L}^{*}(\phi^{-})$ to be defined through its cash flows. Inserting $\phi_{t}=\phi_{t}^{+}-\phi_{t}^{-}$ into the formal portfolio conversion equation (2) for $\phi$ at date $t\in D$ and rearranging we get a portfolio conversion equation for the non-negative strategy $\phi^{+}$:
@@
In view of this, we define the cash flows for the liability $\mathcal{L}^{*}\equiv\mathcal{L}^{*}(\phi^{-})$ at dates $t$ to be the cash outflow $X_{t}^{\mathcal{L}^{*}}=\phi_{t}^{-}\cdot S_{t}+\tilde{Z}^{\phi^{-}}_{t}$ above from closing out at date $t$ a short position in the portfolio $\phi_{t}^{-}$ and the cash inflows $\tilde{Z}_{t}^{\mathcal{L}^{*}}=\phi_{\gamma(t)}^{-}\cdot S_{t}$ above from taking on a new short position in the portfolio $\phi_{\gamma(t)}^{-}$ provided that $t<t_{max}$.
@@
Assume that portfolios containing negative units of tradables are available for production with close out. A strategy $\phi=\phi^{+}-\phi^{-}$ on $D\equiv D[t_{min},t_{max}]$ with cash flows $\tilde{Z}_{t}$ incoming to and cash flows $X_{t}$ outgoing of the strategy is defined as the non-negative strategy $\phi^{+}$ with cash flows $\tilde{Z}_{t}$ and $X_{t}$ together with the liability $\mathcal{L}^{*}\equiv\mathcal{L}^{*}(\phi^{-})$ defined by the cash inflows and outflows for any $t\in D$,
@@
Restrictions. In certain applications, strategies may have to be restricted e.g. to investing only in a subset of tradables, i.e. $\phi_{t}^{k}=0$ for one or several $k\in\{0,1,\ldots d\}$.
@@
For a linear subspace $\mathcal{R}$ of $\mathbb{R}^{d+1}$ and a strategy $\phi$ on $D\equiv D[t_{min},t_{max}]$, we define, by slight abuse of notation, $\phi\in\mathcal{R}$ iff $\phi_{t}\in\mathcal{R}$ for any $t\in D$, $\phi\in\mathcal{R}^{\geq 0}$ iff $0\leq\phi_{t}\in\mathcal{R}$ for any $t\in D$, and ϕitalic-$\phi\in\mathcal{R}^{\prime}$ iff $\phi_{t}\in\mathcal{R}$ and $v_{t}(\phi)\geq 0$ for any $t\in D$. (Note: $\phi\in\mathcal{R}\Rightarrow\phi^{\pm}\in\mathcal{R}$.)